OpenStudy (kimo_indie):

An equation in the system is y = x + 1. What is the other equation in the system if the solution is only at (0, 1)? a) x − y = 1 b) 2x + 2y = 4 c) x + y = 1 d) x + 2y = 2

OpenStudy (trojanpoem):

Plug in that value, the equation which verifies it is the correct choice: a) L.H.S = 0 - 1 = -1 b) L.H.S = 0 +2 = 2 c) 0 + 1 = 1 = R.H.S d) 0 + 2 = 2 = R.H.S